OpenStudy (anonymous):

(4x- 3)^3 please expandand show workings !

OpenStudy (raden):

(4x- 3)^3 = (4x-3)(4x-3)(4x-3)

OpenStudy (anonymous):

yes i know how to do that i just dont know how to get the answer

OpenStudy (raden):

u can find the multiplication if 2 terms, first (4x-3)(4x-3) = ...

OpenStudy (mertsj):

\[(a-b)^3=a^3-3a^2b+3ab^2-b^3\]

OpenStudy (mertsj):

Use that and replace each a with 4x and each with 3